OpenStudy (anonymous):

what is the product of 0.8 and 0.3

OpenStudy (anonymous):

.24 if you mean .8(.3)

OpenStudy (e.cociuba):

To find the product of 0.8 and 0.3, all you have to do is multiply. So, tell me. What do you get when you multiply 0.8 x 0.3= ?

OpenStudy (zzr0ck3r):

\[.8 = \frac{8}{10}\\.3 = \frac{3}{10}\\.8*.3=\frac{8}{10}*\frac{3}{10}=\frac{24}{100}=.24\]

OpenStudy (zzr0ck3r):

this is how I like to do them....